i need to know how to convert 19/20 into a percentage and decimal in algebra form
Answer by stanbon(75887)

i need to know how to convert 19/20 into a percentage and decimal in algebra form
----------------
Percentage means "per one-100", so use a proportion.
-----------------------
x/100 = 19/20
x = 100(19/20) = 5*19 = 95
-----------------------------
19/20 = 95%
================================
Cheers,
Stan H.